# Deliberate Planning in Language Models with Symbolic Representation

### v2

@@ 3.1 Experimental Setup
We evaluate SymPlanner on PlanBench (Valmeekam et al., 2023), a symbolic planning benchmark where a robotic agent manipulates blocks on a table to achieve a target configuration. PlanBench is particularly suitable for our study because it requires strict adherence to symbolic constraints while supporting long-horizon reasoning. The action space consists of four operations (Pickup, Putdown, Stack, and Unstack), which are represented and parsed by language models (Yang et al., 2023).
@@
We compare SymPlanner with three widely used LM-based reasoning paradigms: chain-of-thought (CoT), tree-of-thought (ToT), and reasoning-as-planning (RAP), implemented via the LLM Reasoners library (Hao et al., 2024). For CoT, we generate independent reasoning chains with a fixed temperature of 0. ToT is configured with a generation limit of $4$, a breadth of $4$, and a search depth of $6$, resulting in $32$ total rollouts. RAP is set with a depth limit of $6$, cumulative reward aggregation, mean-based Q-value estimation, and uses Monte Carlo tree search (MCTS) with symbolic feedback. Verification for all methods is performed with a symbolic validator (Howey et al., 2004).
@@
Table 1: Overall accuracy (%) of different methods on PlanBench across varying step counts and base models. Plan validity is checked using symbolic evaluation. Bold numbers indicate the best result for each model.
@@ 3.2 Main Results
Table 1 reports symbolic planning accuracy across different step counts for GPT-4o-mini, GPT-4o, and GPT-4.1. Accuracy is measured as the percentage of generated plans that achieve the goal state under symbolic validation, ensuring that only logically consistent and executable action sequences are counted as correct.
@@
SymPlanner consistently achieves the highest overall accuracy, and the gains become more pronounced with larger base models. For example, SymPlanner reaches $50.0\%$ accuracy with GPT-4o and $54.2\%$ with GPT-4.1, compared to only $21.6\%$ with GPT-4o-mini. This scaling trend highlights two important points. First, larger models provide stronger priors for generating semantically meaningful actions, which SymPlanner can further refine through symbolic validation. Second, the improvement gap between small and large models suggests that SymPlanner is capable of exploiting model capacity more effectively than natural language–only baselines, which often plateau or collapse as task difficulty increases. In other words, the framework not only scales with model size but also amplifies the benefits of stronger base LMs.
@@
Among the baselines, few-shot CoT outperforms zero-shot CoT, confirming the benefit of demonstration examples. However, even with four-shot prompting, CoT remains far below SymPlanner, particularly at higher step counts, indicating that prompting improvements alone are insufficient for overcoming error accumulation in long-horizon reasoning. ToT, while more structured than CoT, quickly degrades beyond four steps, reflecting its reliance on breadth-first exploration without an explicit mechanism for grounding or correction. RAP achieves strong results at very short horizons (e.g., $95.0\%$ at two steps with GPT-4.1), but its accuracy collapses as depth increases, underscoring its reliance on local step-by-step optimization without global consistency. Collectively, these results demonstrate that current reasoning paradigms lack the mechanisms to effectively constrain or recover from invalid trajectories, making them fragile for complex planning tasks.
@@
In contrast, SymPlanner sustains competitive performance across long horizons, maintaining $55.0\%$ and $50.0\%$ accuracy at eight steps for GPT-4o and GPT-4.1, respectively. The stability of accuracy across longer step counts suggests that iterative correction successfully mitigates cascading errors, while contrastive ranking provides reliable selection among competing partial plans. This robustness contrasts sharply with the steep declines observed in baselines, which often fail once reasoning exceeds a handful of steps. Overall, these findings demonstrate that SymPlanner not only improves single-step decision quality but also maintains coherence over extended trajectories, thereby enabling reliable execution in domains that require structured, multi-step planning under strict symbolic constraints. The combination of scalability, robustness, and symbolic grounding positions SymPlanner as a promising framework for bridging natural language reasoning and formal planning systems.
@@ 3.3 Ablation Study
To assess the contribution of each component in SymPlanner, we conduct an ablation study on GPT-4.1 (Table 2). Accuracy is measured as the proportion of generated plans that achieve the goal under symbolic validation. The complete method achieves the highest overall accuracy of $54.2\%$, confirming that all three mechanisms are beneficial. 1) Symbolic representation. Removing the symbolic world model reduces accuracy to $46.7\%$. While short-horizon tasks remain solvable, longer trajectories quickly degrade, showing that explicit state tracking is essential for consistency. 2) Iterative correction. Ablating IC yields the sharpest decline, with accuracy falling to $20.8\%$. This confirms IC’s role in preventing single-step errors from compounding, making it the most critical component for long-horizon stability. 3) Contrastive ranking. Without CR, accuracy drops to $27.5\%$, particularly at intermediate horizons. CR enables finer discrimination between superficially similar plans, guiding the search toward coherent trajectories. Overall, these results demonstrate that symbolic grounding, correction, and ranking target complementary failure modes and are all necessary for robust, scalable planning.
@@ 4 Related Work
Enhance LMs with test-time compute.
@@
Enhance LMs with symbolic representation.
Integrating symbolic reasoning with LMs has emerged as a promising direction for improving robustness and interpretability. Symbolic representations provide structured state descriptions, external knowledge interfaces, and mechanisms for verifying action validity (Liu et al., 2024; Ni et al., 2024; Xiong et al., 2024; Yu et al., 2024; He et al., 2023, 2024). Within this paradigm, policy strategies can be organized around goal alignment, which evaluates actions by their proximity to the goal, and effect alignment, which validates consistency of predicted outcomes with symbolic rules (Kirk et al., 2024). SymPlanner follows the effect-alignment perspective by validating actions through a symbolic simulator, while extending prior work with iterative correction to repair invalid actions and contrastive ranking to select coherent long-horizon plans.

### v3

@@ 3.1 Experimental Setup
We evaluate SymPlanner on PlanBench (Valmeekam et al., 2023a), a symbolic planning benchmark where a robotic agent manipulates blocks on a table to achieve a target configuration. PlanBench is particularly suitable for our study because it requires strict adherence to symbolic constraints while supporting long-horizon reasoning. The action space consists of four operations (Pickup, Putdown, Stack, and Unstack), which are represented and parsed by language models (Yang et al., 2023).
@@
We compare SymPlanner with three widely used LLM-based reasoning paradigms: chain-of-thought (CoT), tree-of-thought (ToT), and reasoning-as-planning (RAP), implemented via the LLM Reasoners library (Hao et al., 2024). For CoT, we generate independent reasoning chains with a fixed temperature of 0. ToT is configured with a generation limit of $4$, a breadth of $4$, and a search depth of $6$, resulting in $32$ total rollouts. RAP is set with a depth limit of $6$, cumulative reward aggregation, mean-based Q-value estimation, and uses Monte Carlo tree search (MCTS) with symbolic feedback. Verification for all methods is performed with a symbolic validator (Howey et al., 2004).
@@
Table 1: Success rate (%) on PlanBench by optimal plan length. A plan counts as a success only if the simulator validates all steps and the final state entails the goal. Mean over 20 instances per length; 95% CIs in parentheses. Plan validity is checked using symbolic evaluation. Bold numbers indicate the best result for each model.
@@ 3.2 Main Results
Table 1 reports symbolic planning success rate across different step counts for GPT-4o-mini, GPT-4o, and GPT-4.1. Success rate is measured as the percentage of generated plans that achieve the goal state under symbolic validation, ensuring that only logically consistent and executable action sequences are counted as correct.
@@
SymPlanner consistently achieves the highest overall success rate, and the gains become more pronounced with larger base models. For example, SymPlanner reaches $50.0\%$ success rate with GPT-4o and $54.2\%$ with GPT-4.1, compared to only $21.6\%$ with GPT-4o-mini. This scaling trend highlights two important points. First, larger models provide stronger priors for generating semantically meaningful actions, which SymPlanner can further refine through symbolic validation. Second, the improvement gap between small and large models suggests that SymPlanner is capable of exploiting model capacity more effectively than natural language–only baselines, which often plateau or collapse as task difficulty increases. In other words, the framework not only scales with model size but also amplifies the benefits of stronger base LLMs.
@@
Among the baselines, few-shot CoT outperforms zero-shot CoT, confirming the benefit of demonstration examples. However, even with four-shot prompting, CoT remains far below SymPlanner, particularly at higher step counts, indicating that prompting improvements alone are insufficient for overcoming error accumulation in long-horizon reasoning. ToT, while more structured than CoT, quickly degrades beyond four steps, reflecting its reliance on breadth-first exploration without an explicit mechanism for grounding or correction. RAP achieves strong results at very short horizons (e.g., $95.0\%$ at two steps with GPT-4.1), but its success rate collapses as depth increases, underscoring its reliance on local step-by-step optimization without global consistency. Collectively, these results demonstrate that current reasoning paradigms lack the mechanisms to effectively constrain or recover from invalid trajectories, making them fragile for complex planning tasks.
@@
In contrast, SymPlanner sustains competitive performance across long horizons, maintaining $55.0\%$ and $50.0\%$ success rate at eight steps for GPT-4o and GPT-4.1, respectively. The stability of success rate across longer step counts suggests that iterative correction successfully mitigates cascading errors, while contrastive ranking provides reliable selection among competing partial plans. This robustness contrasts sharply with the steep declines observed in baselines, which often fail once reasoning exceeds a handful of steps. Overall, these findings demonstrate that SymPlanner not only improves single-step decision quality but also maintains coherence over extended trajectories, thereby enabling reliable execution in domains that require structured, multi-step planning under strict symbolic constraints. The combination of scalability, robustness, and symbolic grounding positions SymPlanner as a promising framework for bridging natural language reasoning and formal planning systems.
@@ 3.3 Ablation Study
To assess the contribution of each component in SymPlanner, we conduct an ablation study on GPT-4.1 (Table 2). Success rate is measured as the proportion of generated plans that achieve the goal under symbolic validation. The complete method achieves the highest overall success rate of $54.2\%$, confirming that all three mechanisms are beneficial. 1) Symbolic representation. Removing the symbolic world model reduces success rate to $46.7\%$. While short-horizon tasks remain solvable, longer trajectories quickly degrade, showing that explicit state tracking is essential for consistency. 2) Iterative correction. Ablating IC yields the sharpest decline, with success rate falling to $20.8\%$. This confirms IC’s role in preventing single-step errors from compounding, making it the most critical component for long-horizon stability. 3) Contrastive ranking. Without CR, success rate drops to $27.5\%$, particularly at intermediate horizons. CR enables finer discrimination between superficially similar plans, guiding the search toward coherent trajectories. Overall, these results demonstrate that symbolic grounding, correction, and ranking target complementary failure modes and are all necessary for robust, scalable planning.
@@ 4 Related Work
Enhance LLMs with test-time compute.
@@
Enhance LLMs with symbolic representation.
Integrating symbolic reasoning with LLMs has emerged as a promising direction for improving robustness and interpretability. Symbolic representations provide structured state descriptions, external knowledge interfaces, and mechanisms for verifying action validity (Liu et al., 2024; Ni et al., 2024; Xiong et al., 2024a; Yu et al., 2024; He et al., 2023, 2024; Wang et al., 2025; Valmeekam et al., 2023b). Within this paradigm, policy strategies can be organized around goal alignment, which evaluates actions by their proximity to the goal, and effect alignment, which validates consistency of predicted outcomes with symbolic rules (Kirk et al., 2024). SymPlanner follows the effect-alignment perspective by validating actions through a symbolic simulator, while extending prior work with iterative correction to repair invalid actions and contrastive ranking to select coherent long-horizon plans.